OpenStudy (anonymous):

My answer is D Which of the following statements is true? (1 points) Answer An exponential function is the inverse of a logarithmic function. The base of the logarithmic function f(x) = log5x is unknown. The range of the logarithmic function f(x) = log5x is all real numbers less than zero. The domain of the logarithmic function f(x) = log5x is all real numbers.

OpenStudy (anonymous):

The domain of the logarithmic function f(x) = log5x is all real numbers.

OpenStudy (anonymous):

is this for a test

jimthompson5910 (jim_thompson5910):

can you plug in a negative number into f(x) = log(5, x) ?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

you can plug in x = -2 into f(x) = log(5, x) ?

OpenStudy (anonymous):

oh wait, no you cant

OpenStudy (anonymous):

it has to be a positive, right?

jimthompson5910 (jim_thompson5910):

yes because it's only defined for positive numbers

jimthompson5910 (jim_thompson5910):

exactly

OpenStudy (anonymous):

okay

jimthompson5910 (jim_thompson5910):

so that rules out D

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

omg C is true!!!!!! C is true!!!!!!!!!!!!!1

jimthompson5910 (jim_thompson5910):

is it possible for the output of f(x) = log(5, x) to be a positive number?

OpenStudy (anonymous):

Yeah of course

jimthompson5910 (jim_thompson5910):

so that rules out C

jimthompson5910 (jim_thompson5910):

because choice C says that you only get negative numbers as outputs

OpenStudy (anonymous):

Oh, I didn't see that okay umm let me keep checking

OpenStudy (anonymous):

it's def not b

jimthompson5910 (jim_thompson5910):

yes because the base is known (looks like 5)

jimthompson5910 (jim_thompson5910):

so it's not unknown

OpenStudy (anonymous):

so its A!!!!!!!!!!!!!!!!!!!!!!!!!!

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

and it turns out that exponential functions are indeed inverses of logarithmic functions

OpenStudy (anonymous):

yay while we did that, i finished twelve other questions on my own!!!!!!!!!!!!!

jimthompson5910 (jim_thompson5910):

lol good for you, congrats on that

OpenStudy (anonymous):

thank you for your help.

jimthompson5910 (jim_thompson5910):

yw